# Per-antenna power constraints: constructing Pareto-optimal precoders with cubic complexity under non-negligible noise conditions

### v1

@@ IV-C Pareto-optimality verification
Let us now numerically check whether the developed algorithm does indeed parameterize the full Pareto boundary. Let us invoke the slow interior-point QoS solver coupled with a line-search as a reference and compare the resulting surfaces: such a comparison is provided in Figure 6 in the Appendix. The surfaces on the figure are comprised using less points ($7.8\times 10^{3}$), for two reasons:
@@
  • In order to show parameterization differences visually: since the reference computes an intersection of an arbitrary ray with the Pareto boundary, we have the full control of the selected rays. Hence, we can ’emit’ the rays in perfect circles seen on the reference graph clearly. The Lagrange weigths $\lambda$, however, parameterize the surface in a nonlinear way - while the gereral monotonicity between $\lambda_{k}$ and $\mathrm{SINR}_{k}$ holds, the exact location on the boundary cannot be determined in advance based on $\lambda$.
In three dimensions, we can compare surfaces visually, but we should also check if the developed precoding algorithm accurately arrives on the Pareto boundary in higher dimensions. To do so we devised the following experiment. Assume we fix the problem size and power/noise parameters $m_{tx},m_{ue},\omega,\beta$, but vary the channels $H\in\mathbb{C}^{m_{tx}\times m_{ue}}$ by generating elementwise i.i.d. zero-mean random Gaussian matrices. Assume we also randomly sample $\lambda$ with a uniform distribution in $[0,1]$ and then scale those values towards a sum of 1. Assume for each sample we compute, as before,
@@
for small $\upsilon>0$ that is comparable to $\delta$ - in order to estimate how close $P_{sample}$ is to the achievable boundary. The results are provided in Table II. It can be seen that the algorithm tolerance $\delta$ reflects how close the generated precoder is to the Pareto boundary quite well. As reflected in the table, none of the sample precoders could be line-improved by a factor of $1+10\delta$, which gives a strong empirical evidence that the developed algorithm can be seen as a Pareto-boundary parameterization at both larger and smaller problem sizes.
@@ IV-D Varying ${\cal N}=\omega^{2}$ values: correctness of Proposition 1
The condition (10) of the proved full power necessity Proposition 1 can be seen as a lower bound on noise values ${\cal N}_{k}=\omega^{2}$ if the remaining problem parameters (channel $H$, power constraints $\beta$) are fixed [^7] Note that varying noise variances $\omega_{k}^{2}$ is equivalent to varying the power constraint values $\beta_{k}$ by the definition of SINR as in (1).. In previous experiments, $\omega$ was set to a high value of $1$, and experimentally every single run of Algorithm 1, coupled with $\mu$-refinement of Subsection III-A, has converged.
Let us study how the theory and practice break down if $\omega\rightarrow 0$ on the same toy channel (31). Figure 5 in the Appendix shows how the results of the developed algorithm change for reduced $\omega$. Note that the surfaces slowly transform into a cubic shape. The graph that corresponds to $\omega=0.04$ contains ’ruptures’ - areas in which one of elements $\mu_{i}$ converges to zero.
Using the ’rupture’ locations we managed to extract a precoder $P_{example}$ presented in Table V in the Appendix. The reader can verify that $P_{example}$ is Pareto-optimal, but doesn’t utilize full power. Moreover, it was verified that Lemma 2 indeed holds and $H^{*}P_{example}\mathrm{diag}(v)$ has an eigenvalue of $1$. That confirms the main ideas of our theoretical analysis, but it doesn’t mean that the developed precoding algorithm is unusable for lower noise parameters $\omega$. The computationally-efficient version of Algorithm 1 doesn’t accept exact zero values of $\mu$, but if a thresholding is implemented to guarantee $\mu_{j}\geq\mu_{\min}>0$, the developed algorithm still works - we did not do that on Figure 5c for demonstration purposes. Moreover, the most valuable points of the performance region seem unaffected by the ’ruptures’ anyway.
@@ IV-E Varying sizes and ${\cal N}=\omega^{2}$ values: gain asymptotics
Let us study the gains provided by the developed algorithm in higher dimensions $m_{tx}>m_{ue}>3$. We will not be using a fixed toy channel example anymore, so let us define ground rules we will follow while comparing experimental results with different size, channel model and norm specifications.
@@
Empirically, these rules guarantee that with a fixed $\chi$ individual throughput values (32) are approximately constant regardless of $m_{tx},m_{ue}$, at least on randomly generated channels. The channels we will be considering first will be constructed using the SVD representation (27) with random othogonal bases U,VU,V [^8] U,VU,V were constructed as the row/column orthogonal bases of a random Gaussian matrix in experiments. and predefined singular value decay law in one of the three forms:
@@
If $m_{ue}>3$, we are unable to compare whole surfaces, and we have to restrain to comparing individual precoder performance; equivalently, we will be comparing individual points on the surfaces shown in previous experiments. The point selection strategies will very roughly optimize mean/total throughput (4):
@@
    • Water-filling. The weights $\kappa$ are selected in such a way [^9] Water-filling weights that optimize ZF throughput are used for the SLNR baseline aswell. that $P_{\mathrm{ZF}}D_{\kappa}$ is optimal in terms of throughput (4) under Frobenius power constraint (3) [6].
@@
Let us then study the gain dependencies on the problem sizes $m_{tx},m_{ue}$. The Figures 4c, 4d, 4e in the Appendix present example how the values (34) behave depending on the problem sizes $m_{tx},m_{ue}$ for various channel singular value decay laws. Graphically, while the gain factor remains strictly above 1, it can be seen that even the monotonicity of the gain factor dependence on $m_{ue}$ may change for different channel singular value decay laws. We attempted to fit the experimental values with a polynomial model of the form
${\cal G}_{avf}(m_{tx},m_{ue})\leftrightarrow cm_{tx}^{\phi_{tx}}m_{ue}^{\phi_{rx}},$
@@
We have to conclude that the polynomial model power fittings $\phi_{tx},\phi_{ue}$ largely depend on the channel conditioning.
@@ IV-F Experiments on simulated channel
While random Gaussian channels are easier to generate and analyze, wireless communication algorithms are commonly tested on software-simulated channels, designed to closer represent practical application requirements. However, we claim that no qualitative changes appear if we plug in complex-valued simulated values built using the known QuadRiga software. In order to mimic the previous experiment, we regenerated 40 channel samples using the ’BERLIN_UMa_NLoS’ model for a large variety of size specifications [^11] Varied $m_{ue}$ values were simulated by selecting user subsets; varied $m_{tx}$ values corresponded to separate QuadRiGa runs..
The results are provided on Figures 4f, 4b, 4a along with the randomly-generated channel experiments for comparison. One can see that the SINR multiplicative gain is approximately constant (each individual SINR value commonly doubles), which leads to constant additive gain in the mean throughput values again.
@@ IV-G Global optimization
  • It is possible to utilize a black-box derivative-free optimizer over $\lambda$. There is a variety of options, ranging from simpler Nelder-Mead [23] algorithm up to a sophisticated Tensor-Train optimization technique [4], that, assuming smoothness, approximates the target function value grid with a high-dimensional low-rank tensor while computing only a small fraction of its elements. With the Tensor-Train optimizer, we have managed to obtain a global optimum of (4) for sizes $H\in\mathbb{C}^{192\times 24}$ in $15$ minutes of computing time on a desktop Intel Core i9-7900X processor. While this option is surely not realtime-applicable, it is significantly faster than the QoS solver coupled with the Branch-reduce-bound method, which we managed to converge with at sizes $m_{tx}<10,m_{ue}<10$ only.
@@
  • If the total/mean throughput is requested, then equal $\lambda_{k}=\frac{1}{m_{ue}}$ happens to be close to global optimum parameterization often, so maybe exact optimization is not so necessary? In applications, however, the global optimization functional commonly applies weights to throughputs to simulate ’proportional fairness’. One could study how these weights affect the ’average’ close-to-optimal $\lambda_{k}$.
@@ V Conclusion
It was shown experimentally that even two iterations (’starter’ step and one $\mu$ update) are sufficient to build a family of precoders than noticeably outperform ZF/SLNR baselines, while 10 iterations are sufficient to converge to the Pareto boundary very accurately - so that the obtained SINR combination cannot be jointly improved by a factor of $1+10^{-3}$.
Our numerical study shows that the per-user relative SINR gain the Pareto-optimal precoder offer against ZF/SLNR baselines increases with $m_{tx}$ and decreases with $m_{ue}$ for well-conditioned channels, but changes mildly with the problem sizes. Very roughly that means that the per-user throughputs, which are in logarithmic dependence over the SINR values, can be increased by a constant regardless of the initial throughput value - but this constant is larger for ill-conditioned channels.
To sum up, the developed algorithm, to our knowledge, brings the most efficient way to construct Pareto-optimal precoders under per-antenna power constraints, but still at a constant-factor-higher price as compared to the baselines of ZF/SLNR. It can be effectively used either ’as is’ or as an auxiliary tool to perform global throughput optimization over the feasible SINR region boundary. Whether the gain offered by Pareto-optimal or globally optimal precoders with respect to baselines is worth the additional cost or not - is, of course, dependent on the application.
[Numerical tables and figures] Due to their large double-column sizes, we arranged some of the numerical result Figures and Tables as the Appendix at the end of the paper.

### v2

@@ IV-C Pareto-optimality verification
Let us now numerically check whether the developed algorithm does indeed parameterize the full Pareto boundary. Let us invoke the slow interior-point QoS solver coupled with a line-search as a reference and compare the resulting surfaces: such a comparison is provided in Figure 6. The surfaces on the figure are comprised using less points ($7.8\times 10^{3}$), for two reasons:
@@
  • In order to explain parameterization differences visually: the developed precoding algorithm depends on the Lagrange multipliers $\lambda$ in a nonlinear way. While the gereral monotonicity between $\lambda_{k}$ and $\mathrm{SINR}_{k}$ commonly holds, the exact location on the boundary cannot be determined in advance based on underlying $\lambda$.
To verify that the developed algorithm converges to Pareto-optimal points in higher dimensions, we devised the following experiment. Assume we fix the problem size and power/noise parameters $m_{tx},m_{ue},\omega,\beta$, but vary the channels $H\in\mathbb{C}^{m_{tx}\times m_{ue}}$ by generating elementwise i.i.d. zero-mean random Gaussian matrices. Assume we also randomly sample $\lambda$ with a uniform distribution in $[0,1]$ and then scale those values towards a sum of 1. Assume for each sample we compute, as before,
@@
for small $\upsilon>0$ that is comparable to $\delta$ - in order to estimate how close $P_{sample}$ is to the achievable boundary. The results are provided in Table II. It can be seen that the algorithm tolerance $\delta$ reflects how close the generated precoder is to the Pareto boundary quite accurately.
@@ IV-D Varying ${\cal N}=\omega^{2}$ values: correctness of Proposition 1
Let us study how the theory and practice break down if $\omega\rightarrow 0$ on the same toy channel of Figure 1d - which would break the assumption (8) our theoretical analysis of Proposition 1. Figure 4 shows how the results of the developed algorithm change for reduced $\omega$. Note that the surfaces slowly transform into a cubic shape.
The graph that corresponds to $\omega=0.04$ contains ’ruptures’ - areas in which one of elements $\mu_{i}$ converges to zero. Using the ’rupture’ locations we managed to extract a precoder $P_{example}$ presented in Table V in the Supplementary materials. The reader can verify that $P_{example}$ is Pareto-optimal, but doesn’t utilize full power. Moreover, it was verified that Lemma 2 indeed holds and $H^{*}P_{example}\mathrm{diag}(v)$ has an eigenvalue of $1$. That confirms the main ideas of our theoretical analysis, but it doesn’t mean that the developed precoding algorithm is unusable for lower noise parameters $\omega$. The computationally-efficient version of Algorithm 1 doesn’t accept exact zero values of $\mu$, but if a thresholding is implemented to guarantee $\mu_{j}\geq\mu_{\min}>0$, the developed algorithm still works - we did not do that on Figure 4c for demonstration purposes. Moreover, the most valuable points of the performance region seem unaffected by the ’ruptures’ anyway.
(a) Parametric precoder, higher $\omega=1$
(b) Parametric precoder, medium $\omega=0.2$
(c) Parametric precoder, lower $\omega=0.04$
(d) SLNR reference, higher $\omega=1$
(e) SLNR reference, medium $\omega=0.2$
(f) SLNR reference, lower $\omega=0.04$
Figure 4: Varying the SINR noise level value $\omega$ for the three-dimensional toy channel $H\in\mathbb{R}^{8\times 3}$ given by Figure 1d for the developed $P(\lambda,\mu(\lambda))$ and SLNR reference. The graph slowly transforms into a cubic shape as $\omega\rightarrow 0$.
TABLE II: Success rate of improving each SINR value by a factor $\upsilon$, 100 runs of developed algorithm with a threshold $\delta$ on random Gaussian channels
$P\in\mathbb{C}^{24\times 8},\delta=10^{-4}$ 0 % 6 % 47 % 82 % 100 %
$P\in\mathbb{C}^{192\times 24},\delta=10^{-4}$ 0 % 9 % 28 % 44 % 95 %
@@ IV-E Varying sizes and ${\cal N}=\omega^{2}$ values: gain asymptotics
Let us study the gains provided by the developed algorithm in higher dimensions $m_{tx}>m_{ue}>3$. Let us define ground rules we will follow while comparing experimental results with different size, channel model and norm specifications.
@@
Empirically, these rules guarantee that with a fixed $\chi$ individual throughput values $\log(1+\mathrm{SINR}_{k})$ are approximately constant regardless of $m_{tx},m_{ue}$ - at least on randomly generated channels. The channels we will now be considering will be constructed using the SVD representation (25) with random othogonal bases U,VU,V [^9] U,VU,V were constructed as the row/column orthogonal bases of a random Gaussian matrix in experiments. and predefined singular value decay law in one of the three forms:
@@
If $m_{ue}>3$, we are unable to compare whole surfaces, and we have to restrain to comparing individual points on the surfaces shown in previous experiments. The point selection strategies will roughly optimize mean/total throughput (4):
@@
    • Water-filling. The weights $\kappa$ are selected in such a way [^10] The water-filling weights that optimize ZF throughput are used for the SLNR baseline aswell, disregarding the nonzero interferences ${\cal I}_{k}$ of SLNR. that $P_{\mathrm{ZF}}D_{\kappa}$ is optimal in terms of throughput (4) under Frobenius power constraint (3) [6].
@@
Let us then study the gain dependencies on the problem sizes $m_{tx},m_{ue}$. The Figures 5b, 5c, 5e present example how the values (30) behave depending on the problem sizes $m_{tx},m_{ue}$ for various channel singular value decay laws. We attempted to fit the experimental values with a polynomial model of the form
${\cal G}_{avg}(m_{tx},m_{ue})\leftrightarrow cm_{tx}^{\phi_{tx}}m_{ue}^{\phi_{rx}},$
@@
Note that even the sign of $\phi_{ue}$ may change - we have to conclude that the polynomial power fittings $\phi_{tx},\phi_{ue}$ of the gain largely depend on the channel conditioning.
@@ IV-F Experiments on simulated channel
We claim that no qualitative changes appear if we plug in complex-valued simulated values built using the known QuadRiga software. In order to mimic the previous experiment, we regenerated 40 channel samples using the ’BERLIN_UMa_NLoS’ model for a large variety of size specifications [^12] Varied $m_{ue}$ values were simulated by selecting user subsets; varied $m_{tx}$ values corresponded to separate QuadRiGa runs..
The results are provided on Figures 5f, 5d, 5a along with the randomly-generated channel experiments for comparison. One can see that the SINR multiplicative gain is approximately constant (each individual SINR value commonly doubles), which leads to constant additive gain in the mean throughput values again.
@@ IV-G Global optimization
  • It is possible to utilize a black-box derivative-free optimizer over $\lambda$. There is a variety of options, ranging from simpler Nelder-Mead [23] algorithm up to a sophisticated Tensor-Train optimization technique [4]. With the Tensor-Train optimizer, we have managed to obtain a global optimum of (4) for sizes $H\in\mathbb{C}^{192\times 24}$ in $15$ minutes of computing time on a desktop Intel Core i9-7900X processor. While this option is surely not realtime-applicable, it is significantly faster than the QoS solver coupled with the Branch-reduce-bound method, which we managed to converge with at sizes $m_{tx}<10,m_{ue}<10$ only.
@@ V Conclusion
It was shown experimentally that even two iterations (’starter’ step and one $\mu$ update) are sufficient to build a family of precoders that noticeably outperform ZF/SLNR baselines, while 10 iterations are sufficient to converge to the Pareto boundary accurately - so that the obtained SINR combination cannot be jointly improved by a factor of $1+10^{-3}$.
Our numerical study shows that the per-user relative SINR gain the Pareto-optimal precoder offer against ZF/SLNR baselines increases with $m_{tx}$ and decreases with $m_{ue}$ for well-conditioned channels, but changes mildly with the problem sizes. Very roughly that means that the per-user throughputs, in logarithmic dependence over the SINR values, are increased by a constant regardless of the initial throughput value - but this constant is larger for ill-conditioned channels.
To sum up, the developed algorithm, to our knowledge, brings the most efficient way to construct Pareto-optimal precoders under per-antenna power constraints, but still at a constant-factor-higher price as compared to the baselines of ZF/SLNR. It can be effectively used either ’as is’ (by just using equal $\lambda$) or as an auxiliary tool to perform global throughput optimization over the feasible SINR region boundary.
Due to their large double-column size, we arranged Figures 5, 6 as the Appendix at the end of the paper.